A 76 kg adult sits at one end of a 7.6 m long board. His 27 kg child sits on the other end.

A) Where should the pivot be placed so that the board is balanced, ignoring the board's mass?

B) Find the pivot point if the board is uniform and has a mass of 11kg .
Answer #1

a) let the pivot be at a distance x from 76 kg person

m1*x = m2*(L-x)

76*x = 27*(7.6-x)

76*x = 205.2 - 27*x

205.2 = 103*x

x = 1.992 m from 76 kg person

b) let the pivot be at a distance x from 76 kg person

m1*x = m2*(L-x) + m3*(7.6/2 -x)

76*x = 27*(7.6-x) + 11*(3.8 - x)

76*x = 205.2 - 27*x + 41.8 - 11*x

247 = 114*x

x = 2.167 m from 76 kg person